OpenStudy (anonymous):

Which of the following is the solution to the equation radical 5^(8n) = 125^(n + 3)?

OpenStudy (anonymous):

5^8n = 5^3n + 9 8n = 3n + 9 5n = 9 n = 9/5

OpenStudy (anonymous):

n = –9, n = 3, n = 9, and n = –3

OpenStudy (anonymous):

why would it be 9/5?

OpenStudy (anonymous):

If \(5n = 9\) then \(n = 9/5\) o.O

OpenStudy (slaaibak):

I believe he just showed you why.